A quantity of gas has an initial pressure and volume of 0.1 MN/m2 and 0.1 m3. respectively. It is compressed to a final pressure of 1.4 MN/m2 according to the law PV1.26= constant. Determine the final volume of the gas.


1
Expert's answer
2021-10-12T01:58:01-0400
"PV^{1.26}=\text{constant}"

"\dfrac{V_2}{V_1}=(\dfrac{P_1}{P_2})^{1/1.26}"

"V_2=V_1(\dfrac{P_1}{P_2})^{1/1.26}"

"V_2=0.1 m^3\bigg(\dfrac{0.1 MN/m^2}{1.4 MN/m^2}\bigg)^{1/1.26}\approx0.0123\ m^3"

The final volume of the gas is "0.0123" m3..